# Spectral Flow Learning Theory: Finite-Sample Guarantees for Vector-Field Identification

### v3

@@ V-B Observability Inequality and Local Truncation Error (LTE)
Some useful facts we can leverage from the above construction include the following.
@@
$c_{\rm obs}\coloneqq\inf_{\|v\|_{\mathcal{V}}=1}\|\mathsf{B}v\|_{\mathcal{W}}=\sqrt{\lambda_{\min}(\mathsf{B}^{*}\mathsf{B})},$
@@
$\|\mathsf{B}(f-f_{\rho})\|_{\mathcal{W}}^{2}\geq c_{\rm obs}^{2}\|f-f_{\rho}\|_{\mathcal{V}}^{2},\qquad\forall f\in\mathcal{V}.$
@@
Under Assumptions III.1, III.2, and III.3, there exists a constant $C_{\mathrm{LTE},f}>0$, independent of $t$, $h$, and $\zeta$, such that the single-window residual is:
$\|\Delta_{\mathrm{MS}}(f)(x(t),h,\zeta)\|_{2}\leq C_{\mathrm{LTE},f}h^{p+1},$
for all $t\geq 0$, $h\in\boldsymbol{h}$, and $\zeta\in\boldsymbol{\zeta}$. Consequently, in the flow space $L^{2}$ norm,
$\|\Delta_{\mathrm{MS}}(f)\|_{\mathcal{W}}\leq\ C_{\mathrm{LTE},f}\sqrt{\mathbb{E}[H^{2p+2}]}.$ (8)
@@
Let $x(\cdot)$ solve $\dot{x}=x(t)=f(x(t))+g(x(t))u(t)$ on a forward‑invariant compact set $\mathcal{X}$ so that $\sup_{t}\|x^{(p+1)}(t)\|<\infty$. Following the notation in [5], we write the residual componentwise as the variable‑step difference operator:
$\mathcal{L}_{\zeta,h}[x]=x(t{+}h)$
$\quad+\sum_{j=1}^{M-1}\alpha_{j}(\zeta)x(t{-}\tau_{j}(\zeta))-h\sum_{j=0}^{M}\beta_{j}(\zeta)x^{\prime}(t{-}\tau_{j}(\zeta)).$
By Assumption III.2, which provides order-$p$ consistency, Assumption III.1 that guarantees bounded variable‑step coefficients, and Assumption III.3 for smooth ODE, [5, Eq. 5.17] yields:
$\mathcal{L}_{\zeta,h}[x]=\mathcal{O}(h^{p+1})\qquad\text{for smooth }x(\cdot),$
uniformly over admissible $\zeta$. A trivial extension of [5, Lem. 2.2] (constant‑step local error equivalence) to variable steps then implies that the local error at the anchor point is $\mathcal{O}(h^{p+1})$. Identifying $\Delta_{\mathrm{MS}}(f)=\mathcal{L}_{\zeta,h}[x]$ gives the pointwise bound with a constant $C_{\mathrm{LTE},f}$ independent of $(t,h,\zeta)$. This proves the first claim.
For the $\mathcal{W}$ norm, square and integrate over $\rho^{\rm design}_{\mathcal{X},\boldsymbol{h},\boldsymbol{\zeta}}$ to get:
$\|\Delta_{\mathrm{MS}}(f)\|_{\mathcal{W}}^{2}$ $=\int\|\Delta_{\mathrm{MS}}(f)(x,h,\zeta)\|^{2}_{2}\ d\rho^{\rm design}_{\mathcal{X},\boldsymbol{h},\boldsymbol{\zeta}}$
$\leq\ C_{\mathrm{LTE},f}^{2}\int h^{2p+2}d\rho_{\boldsymbol{h}}=C_{\mathrm{LTE},f}^{2}\mathbb{E}[H^{2p+2}].$
Taking the square root on both sides yields (8). ∎
Lemmas 1 and 2 provide stepping stones needed to lift flow-space guarantees to the vector field: a $\mathsf{B}$-observability inequality that controls $\|f-f_{\rho}\|_{\mathcal{V}}$ by $\|\mathsf{B}(f-f_{\rho})\|_{\mathcal{W}}$, and a uniform bound of the vLMM modeling residual via the moment $\mathbb{E}[H^{2p+2}]$. Next, we formalize the flow regression problem, introduce the excess risk $\mathcal{E}(\Phi)$ and the target regression function $\Phi_{\rho}$, and state the flow FS–HP bound obtained by spectral regularization. Combining that bound with the bounded label operator $y_{\alpha}=\mathsf{J}+\mathsf{A}$ and the decomposition $\mathsf{B}(\widehat{f}-f_{\rho})=\Delta_{\rm MS}(\widehat{f})+{\mathsf{Y}}_{\alpha}(\widehat{\Phi}-\Phi_{\rho})+\Delta_{\rm MS}(f_{\rho})$ yields a two-term vector-field bound featuring a statistical rate in $\ell$ and a deterministic vLMM bias in $\mathbb{E}[H^{2p+2}]$, scaled by $c_{\rm obs}$.
V-C From Flow FS-HP Bound to Vector‑Field FS-HP Bound
With flow label $Y\in\mathcal{X}$ and $\rho((x,h),y)=\rho(y\mid x,h)\rho_{\mathcal{X},\boldsymbol{h}}^{\rm design}(x,h)$, the regression function is:
@@ V-C From Flow FS-HP Bound to Vector‑Field FS-HP Bound
For a flow predictor $\Phi\in\mathcal{W}$,
$\mathcal{E}(\Phi)$ $\coloneqq\int_{\mathcal{Z}\times\mathcal{X}}\|\Phi^{h}(x)-y\|^{2}d\rho((x,h),y)$ (9)
$=\|\Phi-\Phi_{\rho}\|_{\mathcal{W}}^{2}+\mathcal{E}(\Phi_{\rho}).$
Thus minimizing $\mathcal{E}$ is equivalent to approximating $\Phi_{\rho}\in\mathcal{W}$ [4, Def. 1]. The choice of the estimated flow $\widehat{\Phi}\coloneqq\Phi_{z,y}^{\lambda_{\ell}}$ is inferred from data ${(z_{i},y_{i})}_{i=1}^{\ell}$ and induced from inductive bias encoded in $\mathcal{H}_{\mathcal{V}}$ through a kernel $k_{\mathcal{V}}$. Therefore, the empirical flow without vLMM modeling error is:
$\widehat{\Phi}\coloneqq\Phi_{z,y}^{\lambda_{\ell}}=\Phi(\ \cdot\ ;f_{z,y}^{\lambda_{\ell}}).$
Notice that the excess risk $\mathcal{E}(\widehat{\Phi})-\mathcal{E}(\Phi_{\rho})=\|\widehat{\Phi}-\Phi_{\rho}\|_{\mathcal{W}}^{2}$ comes directly from (9), and similarly for $\mathcal{E}(\widehat{f})-\mathcal{E}(f_{\rho})$. Under the state design law $\rho_{\mathcal{X}}^{\mathrm{design}}$, flow and vector field excess risk are probabilistically bounded as follows.
Theorem 1 (Flow FS-HP bound).
Suppose $\Phi_{\rho}\in\Omega^{(\mathcal{W})}_{r,R}$ and $r\leq\mu$, which is the qualification $\mathsf{g}_{\lambda}$. Let $\beta=\max\{1,2\mu\}$. For $0<\eta\leq 1$, the following inequality holds with probability at least $1-\eta$:
$\mathcal{E}(\widehat{\Phi})-\mathcal{E}(\Phi_{\rho})\leq\log\frac{4}{\eta}C_{\rm flow}\ell^{-\frac{2r}{2r+\beta}},$ (10)
for some finite positive constant $C_{\rm flow}$.
@@
[4, Thm. 2] applies unchanged to $\mathcal{H}_{\mathcal{W}}\subset\mathcal{W}$. ∎
@@
Suppose $\Phi_{\rho}\in\Omega^{(\mathcal{W})}_{r,R}$ and $r\leq\mu$, which is the qualification $\mathsf{g}_{\lambda}$. For $0<\eta\leq 1$, the following inequality holds with probability at least $1-\eta$:
$\mathcal{E}(\widehat{f})-\mathcal{E}(f_{\rho})\leq\frac{C_{\rm fit}}{c_{\rm obs}^{2}}\log\frac{4}{\eta}\ell^{-\frac{2r}{2r+\beta}}+\frac{C_{\rm bias}}{c_{\rm obs}^{2}}\ \mathbb{E}[H^{2p+2}].$ (11)
for some finite positive constant $c_{\rm obs}$, $C_{\rm fit}$, and $C_{\rm bias}$.
@@
$\|\widehat{f}-f_{\rho}\|_{\mathcal{V}}^{2}$ $\leq\frac{1}{c_{\rm obs}^{2}}\ \|\mathsf{B}(\widehat{f}-f_{\rho})\|_{\mathcal{W}}^{2}$ (12)
$\leq\frac{1}{c_{\rm obs}^{2}}(\|\underbrace{\mathsf{B}\widehat{f}-\mathsf{Y}_{\alpha}\widehat{\Phi}}_{\Delta_{\rm MS}(\widehat{f})}\|_{\mathcal{W}}^{2}\ +\|\mathsf{Y}_{\alpha}(\widehat{\Phi}-\Phi_{\rho})\|_{\mathcal{W}}^{2}$
$\qquad+\|\underbrace{\mathsf{Y}_{\alpha}\Phi_{\rho}-\mathsf{B}f_{\rho}}_{\Delta_{\rm MS}(f_{\rho})}\|_{\mathcal{W}}^{2})$ (13)
$\leq\frac{1}{c_{\rm obs}^{2}}\ \|\mathsf{J}+\mathsf{A}\|^{2}_{\mathcal{L}(\mathcal{W})}\big(\log\frac{4}{\eta}C_{\rm flow}\ell^{-\frac{2r}{2r+\beta}}\big)$
$\qquad+\frac{1}{c_{\rm obs}^{2}}\big(C^{2}_{{\rm LTE},\widehat{f}}+C^{2}_{{\rm LTE},f_{\rho}}\big)\mathbb{E}[H^{2p+2}],$ (14)
@@
First, inequality (12) follows immediately from the Lemma 1. Indeed, by definition $c_{\rm obs}^{2}=\lambda_{\min}(\mathsf{B}^{*}\mathsf{B})$, so for any $v\in\mathcal{V}$, $c_{\rm obs}^{2}\|v\|_{\mathcal{V}}^{2}\leq\|\mathsf{B}v\|_{\mathcal{W}}^{2}$; applying this with $v=\widehat{f}-f_{\rho}$ yields:
$\|\widehat{f}-f_{\rho}\|_{\mathcal{V}}^{2}\leq\frac{1}{c_{\rm obs}^{2}}\|\mathsf{B}(\widehat{f}-f_{\rho})\|_{\mathcal{W}}^{2},$
@@
$\mathsf{B}(\widehat{f}-f_{\rho})=\underbrace{\mathsf{B}\widehat{f}-\mathsf{Y}_{\alpha}\widehat{\Phi}}_{\Delta_{\rm MS}(\widehat{f})}+\underbrace{\mathsf{Y}_{\alpha}(\widehat{\Phi}-\Phi_{\rho})}_{\text{flow mismatch}}+\underbrace{\mathsf{Y}_{\alpha}\Phi_{\rho}-\mathsf{B}f_{\rho}}_{\Delta_{\rm MS}(f_{\rho})},$
@@
Invoking the flow FS–HP bound (Theorem 1) for $\|\widehat{\Phi}-\Phi_{\rho}\|_{\mathcal{W}}^{2}$ then yields the first term appearing in (14).
@@
For the second term on the right-hand side of (14), i.e., the contribution of the two vLMM residuals, we apply Lemma 2 to the first and third terms of (13). This gives $\|\Delta_{\rm MS}(\widehat{f})\|_{\mathcal{W}}^{2}\leq C_{{\rm LTE},\widehat{f}}^{2}\mathbb{E}[H^{2p+2}],$ and $\|\Delta_{\rm MS}(f_{\rho})\|_{\mathcal{W}}^{2}\leq C_{{\rm LTE},f_{\rho}}^{2}\mathbb{E}[H^{2p+2}],$ which together produce the second term in (14).
@@
Finally, note that $\|\mathsf{J}+\mathsf{A}\|_{\mathcal{L}(\mathcal{W})}$ is bounded under Assumption III.1. Therefore we can absorb constants by setting $C_{\rm fit}\geq C_{\rm flow}\|\mathsf{J}+\mathsf{A}\|_{\mathcal{L}(\mathcal{W})}^{2}$ and $C_{\rm bias}\coloneqq C_{{\rm LTE},\widehat{f}}^{2}+C_{{\rm LTE},f_{\rho}}^{2}$. Substituting these into (14) completes the proof of (11). ∎

### v4

@@ V-B Observability Inequality and Local Truncation Error (LTE)
Some useful consequences we can leverage from the above construction include the following.
@@
$c_{\rm obs}(h)\coloneqq\inf_{\|v\|_{\mathcal{V}}=1}\|\mathsf{B}v\|_{\mathcal{W}}=\sqrt{\lambda_{\min}(\mathsf{B}^{*}\mathsf{B})};$
@@
$\|\mathsf{B}(f-f_{\rho})\|_{\mathcal{W}}^{2}\geq c_{\rm obs}(h)^{2}\|f-f_{\rho}\|_{\mathcal{V}}^{2},\qquad\forall f\in\mathcal{V}.$
@@
Under Assumptions III.1, III.2, and III.3, there exists a constant $C_{\mathrm{LTE}}>0$, independent of $t$ and $T$, such that the single-window residual is:
$\|\Delta(f)(x(t);T)\|_{2}\leq C_{\mathrm{LTE}}h^{p+1},$
for all $t\geq 0$, $T\in\mathcal{T}$. Consequently, in the flow space $L^{2}$ norm,
$\|\Delta(f)\|_{\mathcal{W}}\leq\ C_{\mathrm{LTE}}\sqrt{\mathbb{E}[H^{2p+2}]}.$ (8)
@@
Let $x(\cdot)$ solve $\dot{x}=f(x(t))+g(x(t))u(t)$ on a forward‑invariant compact set $\mathcal{X}$ so that $\sup_{t}\|x^{(p+1)}(t)\|<\infty$, where $x^{(p+1)}$ denotes the (p+1)(p{+}1)-order derivate of $x$. Following the notation in [8], we write the residual componentwise as the variable‑step difference operator:
$\mathcal{L}(x;T)=x(t{+}h)$
$\quad+\sum_{j=1}^{M-1}\alpha_{j}(\zeta)x(t{-}\tau_{j}(\zeta))-h\sum_{j=0}^{M}\beta_{j}(\zeta)\dot{x}(t{-}\tau_{j}(\zeta)).$
By Assumption III.2, which provides order-$p$ consistency, Assumption III.1 that guarantees bounded variable‑step coefficients, and Assumption III.3 for a smooth ODE, [8, Eq. 5.17] yields:
$\mathcal{L}(x;T)=\mathcal{O}(h^{p+1}),$
uniformly over admissible $T$ for sufficiently smooth $x(\cdot)$. A trivial extension of the constant‑step local truncation error lemma [8, Lem. 2.2] to variable steps then implies that the local error at each anchor point is $\mathcal{O}(h^{p+1})$. Identifying $\Delta(f)=\mathcal{L}(x;T)$ gives the pointwise bound with a constant $C_{\mathrm{LTE}}$ independent of $(t,T)$. This proves the first claim.
For the $\mathcal{W}$ norm, square and integrate over $\rho^{\rm design}_{\mathcal{X},\mathcal{T}}$ to get:
$\|\Delta(f)\|_{\mathcal{W}}^{2}$ $=\int\|\Delta(f)(x;T)\|^{2}_{2}\ d\rho^{\rm design}_{\mathcal{X},\mathcal{T}}$
$\leq\ C_{\mathrm{LTE}}^{2}\int h^{2p+2}d\rho_{\boldsymbol{h}}=C_{\mathrm{LTE}}^{2}\mathbb{E}[H^{2p+2}].$
Taking the square root of both sides yields (8). ∎
Lemmas 1 and 2 provide stepping stones needed to lift flow-space guarantees to the vector field: a $\mathsf{B}$-observability inequality that controls $\|f-f_{\rho}\|_{\mathcal{V}}$ by $\|\mathsf{B}(f-f_{\rho})\|_{\mathcal{W}}$, and a uniform bound of the vLMM modeling residual via the moment $\mathbb{E}[H^{2p+2}]$. Next, we formalize the flow regression problem, introduce the excess risk $\mathcal{E}(\Phi)$ and the target regression function $\Phi_{\rho}$, and state the flow FS-HP bound obtained by spectral regularization. Combining that bound with the bounded label operator $\mathsf{Y}_{\alpha}=\mathsf{J}+\mathsf{A}$ and the decomposition $\mathsf{B}(\widehat{f}-f_{\rho})=\Delta(\widehat{f})+{\mathsf{Y}}_{\alpha}(\widehat{\Phi}-\Phi_{\rho})+\Delta(f_{\rho})$ yields a two-term vector-field bound featuring a statistical rate in $\ell$ and a deterministic vLMM bias in $\mathbb{E}[H^{2p+2}]$, scaled by $c_{\rm obs}(h)$.
V-C FS-HP Excess Risk Bound for Flow-Learning
We begin by introducing the definition of excess risk, following the formality presented in [11].
Definition 5 (Excess Risk).
The excess risk of a predictor $\Pi$ with respect to its target $\Pi_{\rho}$, i.e., the best possible predictor under a design law $\rho$, is $\mathcal{E}(\Pi)\coloneqq\mathcal{R}(\Pi)-\mathcal{R}(\Pi_{\rho})$, with the expected risk of $\Pi$ and target predictor $\Pi_{\rho}$ given as:
$\mathcal{R}(\Pi)\coloneqq\int_{\mathcal{X}\times\mathcal{Y}}\|\Pi(x)-y\|^{2}d\rho(x,y),\ \Pi_{\rho}(\cdot)\coloneqq\int_{\mathcal{Y}}yd\rho(y\mid\cdot),$
respectively.
With flow label $Y\in\mathcal{X}$ and $\rho((x,T),y)=\rho(y\mid x,h)\rho_{\mathcal{X},\mathcal{T}}^{\rm design}(x,T)$, the target flow predictor is:
@@ V-C FS-HP Excess Risk Bound for Flow-Learning
For any flow predictor $\Phi\in\mathcal{W}$, let $\mathcal{Z}=\mathcal{X}\times\mathcal{T}$:
$\mathcal{R}(\Phi)$ $\coloneqq\int_{\mathcal{Z}\times\mathcal{X}}\|\Phi^{h}(x)-y\|^{2}d\rho((x,T),y)$ (9)
$=\|\Phi-\Phi_{\rho}\|_{\mathcal{W}}^{2}+\mathcal{R}(\Phi_{\rho}).$
Therefore, minimizing $\mathcal{E}(\Pi)$ is equivalent to approximating $\Pi_{\rho}$ with $\Phi$ via minimizing $\|\Pi-\Pi_{\rho}\|$ [11, Def. 1]. The choice of the estimated flow $\widehat{\Phi}\coloneqq\Phi_{\mathcal{D}}^{\lambda_{\ell}}$ is inferred from data $\mathcal{D}={\{z_{i},y_{i}\}}_{i=1}^{\ell}$ and induced from inductive bias encoded in the source condition. Under the state design law $\rho_{\mathcal{X}}^{\mathrm{design}}$, flow excess risk is probabilistically bounded as follows.
Lemma 3 (Flow FS-HP bound).
Suppose $\Phi_{\rho}\in\Omega^{(\mathcal{W})}_{r,R}$ and $r\leq\nu$, which is the qualification $\mathsf{g}_{\lambda}$. For $0<\eta\leq 1$, the following inequality holds with probability at least $1-\eta$:
$\|\Phi-\Phi_{\rho}\|_{\mathcal{W}}\leq\log\frac{4}{\eta}C_{\rm flow}\ell^{-\frac{2r}{2r+\beta}},$ (10)
for some finite positive constant $C_{\rm flow}$ and with $\beta=\max\{1,2\mu\}$.
@@
[11, Thm. 2] applies unchanged to $\mathcal{H}_{\mathcal{W}}\subset\mathcal{W}$. ∎
The rate in Lemma 3 isolates the sample-size $\ell$ dependence from other problem-specific quantities. All other problem- and algorithm-specific quantities are collected into:
$C_{\rm flow}=2C_{1}^{2}+2\gamma_{r}^{2}R^{2},$
with $C_{1}=4\sqrt{2}\kappa M(\sqrt{DB}+\kappa^{5/2}L)$, exactly as in the finite-sample bound in [11, Thm. 2]. Here $\kappa^{2}=\sup_{x}K(x,x)$ bounds the kernel, $M=\sup|x|$ bounds the labels, while $B,D$ and $(L,\mu)$ are the spectral-filter constants defined by $\sup_{0<\sigma\leq\kappa^{2}}|g_{\lambda}(\sigma)|\leq B/\lambda$, $\sup_{0<\sigma\leq\kappa^{2}}|\sigma g_{\lambda}(\sigma)|\leq D$, and $|g_{\lambda}(\sigma)-g_{\lambda}(\sigma^{\prime})|\leq L\lambda^{-\mu}|\sigma-\sigma^{\prime}|$, with $\beta\coloneqq\max\{1,2\mu\}$. The constant $\gamma_{r}$ is the qualification residual at order $r$:
$\sup_{0<\sigma\leq\kappa^{2}}\bigl|1-\sigma g_{\lambda}(\sigma)\bigr|\sigma^{r}\leq\gamma_{r}\lambda^{r},$
which yields the order-$\lambda^{2r}$ approximation term when $\Phi_{\rho}\in\Omega^{(\mathcal{W})}_{r,R}$. All these ingredients combine to give:
$\|\widehat{\Phi}-\Phi_{\rho}\|_{\mathcal{W}}^{2}\leq\log\frac{4}{\eta}\bigl(2C_{1}^{2}+2\gamma_{r}^{2}R^{2}\bigr)\ell^{-\frac{2r}{2r+\beta}}.$
The factor $\log(4/\eta)$ arises from concentration bounds for Hilbert-valued averages. The concentration bounds follow from the Pinelis-Sakhanenko inequality for Hilbert spaces [5, Lemma 8] and can be proved using [14, Theorem 3.3.4].
The flow FS-HP bound is particularly useful when we assume that the target dynamics evolves discretely with a deterministic uniform step schedule, where the step size is irrelevant to the target flow $\Phi_{\rho}$. In SFL, methods targeting such $\Phi_{\rho}$ differ mainly in the hypothesis class $\mathcal{H}_{\mathcal{W}}$, the source condition $\Phi_{\rho}\in\Omega^{(\mathcal{W})}_{r,R_{\mathcal{W}}}=\{\Phi=\mathsf{L}_{\mathcal{W}}^{r}w:\|w\|_{\mathcal{W}}\leq R_{\mathcal{W}}\}$, spectral regularizer $\mathsf{g}_{\lambda}$, and qualification $\nu$. Pipelines are therefore characterized by their choice of $\Omega^{(\mathcal{W})}_{r,R_{\mathcal{W}}}$ and $\mathsf{g}_{\lambda}$, which together govern the rate and stability of the excess risk.
Example V.1 (Methods with Linear/Dictionary Feature Map).
Let $\mathcal{H}_{\mathcal{W}}$ be the vvRKHS induced by a finite dictionary $\psi(x)\in\mathbb{R}^{n}$ with kernel $K(x,x^{\prime})=\psi(x)^{\top}\psi(x^{\prime})$. The population integral operator $\mathsf{L}_{\mathcal{W}}$ acts on this finite-dimensional space. If the model is well specified ($\Phi_{\rho}\in\operatorname{span}\{\psi_{j}\}$), then $\Phi_{\rho}=\mathsf{L}_{\mathcal{W}}^{r}w$ holds for any $r\geq 0$ with $R_{\mathcal{W}}$ determined by the coefficient vector. If misspecified, $r>0$ measures alignment of $\Phi_{\rho}$ with the principal directions of $\mathsf{L}_{\mathcal{W}}$. System identification methods that fall into this category include: ARX and DMDc [10].
On the other hand, Bayesian methods such as Gaussian process state space models (GP-SSM) are also naturally covered by the SFL framework, with a side note that the regularization coefficient $\lambda$ is chosen to be $\sigma^{2}_{y}/\ell$, where $\sigma_{y}$ is the standard deviation of the additive observation noise.
Remark V.1 (TSVD / spectral cut-off).
Spectral cut-off (TSVD), $\mathsf{g}_{\lambda}^{\mathrm{cut}}(\sigma)=\sigma^{-1}\mathbf{1}\{\sigma\geq\lambda\}$, has arbitrarily high qualification but violates the Lipschitz condition used in standard finite-sample proofs [11, 5], so the baseline FS-HP theory does not apply verbatim. Nonetheless, TSVD can be handled within extended spectral-regularization frameworks for non-smooth filters and projection schemes, yielding comparable rates under suitable assumptions; see [6, 2].
Table I provides a more detailed listing of various system identification methods with their corresponding hypothesis class and regularizer $\mathsf{g}_{\lambda}$.
Pipeline Hypothesis in $\mathcal{H}_{\mathcal{W}}$ Filter $\mathsf{g}_{\lambda}$
ARX/ARMAX/BJ Linear/dictionary on $x,u$ Ridge or TSVD
NARX Finite nonlinear dictionary Ridge or TSVD
Subspace ID Past & future Hankel stack TSVD (order choice)
DMD/DMDc Linear in $x,u$ Ridge or TSVD
EDMD/HAVOK Linear in $\psi(x)$ (delays) Ridge or TSVD
KDMD Kernel ridge in $x$ Tikhonov ($\nu{=}1$)
GP-SSM GP mean in RKHS Tikhonov ($\lambda{=}\sigma_{y}^{2}/\ell$)
TABLE I: Classic flow-learning pipelines as spectral-filter estimators in the windowed flow space. Lemma 3 applies verbatim.
V-D FS-HP Excess Risk Bound for Vector-Field Learning
Since discretization bias often matters in practice, we develop an FS-HP excess-risk bound for vector-field learning.
@@ V-D FS-HP Excess Risk Bound for Vector-Field Learning
Suppose $\Phi_{\rho}\in\Omega^{(\mathcal{W})}_{r,R}$ and $r\leq\nu$, which is the qualification $\mathsf{g}_{\lambda}$. For $0<\eta\leq 1$, the following inequality holds with probability at least $1-\eta$:
$\|\widehat{f}-f_{\rho}\|_{\mathcal{V}}^{2}\leq\frac{C_{\rm fit}}{c_{\rm obs}^{2}(h)}\log\frac{4}{\eta}\ell^{-\frac{2r}{2r+\beta}}+\frac{C_{\rm bias}}{c_{\rm obs}^{2}(h)}\ \mathbb{E}[H^{2p+2}].$ (11)
for some finite positive constant $c_{\rm obs}(h)$, $C_{\rm fit}$, and $C_{\rm bias}$.
@@
$\|\widehat{f}-f_{\rho}\|_{\mathcal{V}}^{2}$ $\leq\frac{1}{c_{\rm obs}^{2}(h)}\ \|\mathsf{B}(\widehat{f}-f_{\rho})\|_{\mathcal{W}}^{2}$ (12)
$\leq\frac{1}{c_{\rm obs}^{2}(h)}(\|\underbrace{\mathsf{B}\widehat{f}-\mathsf{Y}_{\alpha}\widehat{\Phi}}_{\Delta(\widehat{f})}\|_{\mathcal{W}}^{2}\ +\|\mathsf{Y}_{\alpha}(\widehat{\Phi}-\Phi_{\rho})\|_{\mathcal{W}}^{2}$
$\qquad+\|\underbrace{\mathsf{Y}_{\alpha}\Phi_{\rho}-\mathsf{B}f_{\rho}}_{\Delta(f_{\rho})}\|_{\mathcal{W}}^{2})$ (13)
$\leq\frac{1}{c_{\rm obs}^{2}(h)}\ \|\mathsf{J}+\mathsf{A}\|^{2}_{\mathcal{L}(\mathcal{W})}\big(\log\frac{4}{\eta}C_{\rm flow}\ell^{-\frac{2r}{2r+\beta}}\big)$
$\qquad+\frac{1}{c_{\rm obs}^{2}(h)}\big(2C^{2}_{{\rm LTE}}\big)\mathbb{E}[H^{2p+2}],$ (14)
@@
First, inequality (12) follows immediately from the Lemma 1. Indeed, by definition, $c_{\rm obs}^{2}(h)=\lambda_{\min}(\mathsf{B}^{*}\mathsf{B})$, so for any $v\in\mathcal{V}$, $c_{\rm obs}^{2}(h)\|v\|_{\mathcal{V}}^{2}\leq\|\mathsf{B}v\|_{\mathcal{W}}^{2}$; applying this with $v=\widehat{f}-f_{\rho}$ yields:
$\|\widehat{f}-f_{\rho}\|_{\mathcal{V}}^{2}\leq\frac{1}{c_{\rm obs}^{2}(h)}\|\mathsf{B}(\widehat{f}-f_{\rho})\|_{\mathcal{W}}^{2},$
@@
$\mathsf{B}(\widehat{f}-f_{\rho})=\underbrace{\mathsf{B}\widehat{f}-\mathsf{Y}_{\alpha}\widehat{\Phi}}_{\Delta(\widehat{f})}+\underbrace{\mathsf{Y}_{\alpha}(\widehat{\Phi}-\Phi_{\rho})}_{\text{flow mismatch}}+\underbrace{\mathsf{Y}_{\alpha}\Phi_{\rho}-\mathsf{B}f_{\rho}}_{\Delta(f_{\rho})},$
@@
Invoking the flow FS-HP bound in Lemma 3 for $\|\widehat{\Phi}-\Phi_{\rho}\|_{\mathcal{W}}^{2}$ then yields the first term appearing in (14).
@@
For the second term on the right-hand side of (14), i.e., the contribution of the two vLMM residuals, we apply Lemma 2 to the first and third terms of (13). This gives $\|\Delta(\widehat{f})\|_{\mathcal{W}}^{2}\leq C_{{\rm LTE}}^{2}\mathbb{E}[H^{2p+2}],$ which produce the second term in (14).
@@
Finally, note that $\|\mathsf{J}+\mathsf{A}\|_{\mathcal{L}(\mathcal{W})}$ is bounded under Assumption III.1. Therefore, we can absorb constants by setting $C_{\rm fit}\geq C_{\rm flow}\|\mathsf{J}+\mathsf{A}\|_{\mathcal{L}(\mathcal{W})}^{2}$ and $C_{\rm bias}\coloneqq 2C_{{\rm LTE}}^{2}$. Substituting these into (14) completes the proof of (11). ∎